What is the volume of hydrogen at a pressure of atm if 200 cm ^ 3 of hydrogen is collected at a pressure of 1,00 atm?
1
Expert's answer
2020-04-18T06:59:22-0400

The final pressure was not specified, that's why it is indicated by: 'x'

p1*V1=p2*V2 ;


V2=p1*V1/p2=1*200/x=___ cm^3